# Exploring the Alignment Landscape: LLMs and Geometric Deep Models in Protein Representation

### v1

@@ Acknowledgment
Appendix A Appendix / Supplemental Material
A.1 Dataset Statistic
Our dataset consists of 20,000 proteins, each with an associated FASTA file and PDB file. Detailed statistics for the dataset are presented in Figure 7, where each pie chart represents a different categorical breakdown of the same dataset. For the "Number of Chains" statistic, the chart shows the count of single-chain proteins versus multiple-chain proteins, with further subdivisions indicating the exact number of chains for multi-chain proteins. For the "Sequence Length" statistic, proteins are categorized by sequence length range. Each section of the pie chart is labeled with the relevant number of chains or sequence length range at the top, with the exact count and corresponding percentage displayed below each label.
A.2 Prompt Used in FASTA Preprocess
@@
Your task is to summarize complex protein sequence data into two
or three sentences that highlight key features such as molecute
type, chains, structural motifs, organism, etc.
@@
A.3 Protein Examples
As shown in Figure 9, we listed several proteins and their corresponding description. The meaning of ‘popular’ and ‘rare’ protein is discussed in Research Question 3: Protein Perspective Analysis.
A.4 GDM Representation Details
A.4.1 GearNet
@@
where $f_{i}$ is the initial feature of node $i$, $\mathcal{N}_{r}(i)$ is the set of neighbors connected to node $i$ via edges of type $r$, $W_{r}$ is a learnable weight matrix for edge type $r$, and $\sigma$ is a ReLU activation function. Batch normalization (BN) is applied for normalization across batches, and a residual connection is added from the previous layer to stabilize training.
This relational message passing allows GearNet to effectively capture both sequential and spatial information in the protein structure, as different edge types are modeled with different convolutional kernels.
@@
where $m^{(l)}_{(i,j,r)}$ represents the message for edge $(i,j,r)$ at layer $l$, $\mathcal{N}^{\prime}_{r^{\prime}}((i,j,r))$ is the set of neighboring edges in $G^{\prime}$, and $W^{\prime}_{r^{\prime}}$ is a learnable weight matrix for edge type $r^{\prime}$. The angular information between edges is used to determine edge types, with edges having smaller angles expected to interact more strongly.
The output from the edge message passing is integrated into the node update step by modifying the aggregation function as follows:
@@
To obtain the final protein representation, the features from all layers are concatenated:
@@
Figure 7: Data Statistic. The pie chart on the left categorizes proteins by the number of chains, while the pie chart on the right categorizes them by sequence length.
A.4.2 GVP
@@
where $148$ is the dimension of the combined scalar and vector features (100 scalar channels and 3 vector channels of size 16, i.e., $3\times 16+100=148$).
@@
Finally, to obtain a fixed-size representation for the entire protein, we apply average pooling across the node dimension:
@@
A.4.3 ScanNet
Parsing the PDB File: First, the PDB file is parsed to extract the protein’s amino acid sequence and atomic point cloud. Each atom is represented as a triplet:
@@ A.4.3 ScanNet
Local Frame Construction: Next, a local reference frame is defined for each heavy atom based on the molecular graph of the protein (i.e., nodes represent atoms and edges represent covalent bonds). For each atom, we select two neighboring atoms to construct a local coordinate system. The center is defined by the atom itself, and the orientation is determined by its neighbors using the following equations:
@@
Atomic and Amino Acid Pooling: ScanNet constructs atomic representations by extracting the local neighborhood of each atom and encoding it with spatio-chemical filters. These filters operate on the atomic coordinates and attributes (such as atom type) and produce an atomic-scale representation. This process is mathematically formalized using spatio-chemical Gaussian filters:
@@
where $G(\mu_{g},\Sigma_{g},x_{k})$ is the Gaussian kernel parameterized by mean $\mu_{g}$ and covariance $\Sigma_{g}$, applied to the local coordinates $x_{k}$ of the neighborhood.
The atomic-scale representations are then pooled at the amino acid level. The pooling step aggregates the atom-wise features into a single representation for each amino acid using a multi-head attention mechanism:
@@
Amino Acid Neighborhood Embedding and Protein Representation: For each amino acid, its local neighborhood is constructed based on the $C_{\alpha}$ atom, sidechain orientation, and backbone orientation. The same spatio-chemical filtering process is applied to obtain an amino acid-wise representation:
@@
where $G(\mu_{g},\Sigma_{g},x_{k})$ is the Gaussian kernel applied to the neighborhood of each amino acid.
@@
A.4.4 GAT
For a graph with $N$ nodes, the input to the GAT layer is a set of node features $\mathbf{h}=\{\mathbf{h}_{1},\mathbf{h}_{2},\dots,\mathbf{h}_{N}\}$, where $\mathbf{h}_{i}\in\mathbb{R}^{F}$ is the feature vector of node $i$, and $F$ is the number of features associated with each atom.
Each node is first transformed linearly using a shared learnable weight matrix $W\in\mathbb{R}^{F^{\prime}\times F}$, yielding: $\mathbf{h}_{i}^{\prime}=W\mathbf{h}_{i}$. Then, the attention mechanism computes attention coefficients for each pair of nodes $i$ and $j$, where $j\in\mathcal{N}_{i}$ is a neighboring node of $i$. The attention coefficients $e_{ij}$ are given by: $e_{ij}=\text{LeakyReLU}\left(\mathbf{a}^{T}\left[W\mathbf{h}_{i}\parallel W\mathbf{h}_{j}\right]\right)$, where $\mathbf{a}\in\mathbb{R}^{2F^{\prime}}$ is a learnable weight vector, and $\parallel$ denotes concatenation. These attention coefficients are then normalized across all neighbors of node $i$ using the softmax function: $\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}.$ The normalized attention coefficients are used to compute the updated feature representation of node $i$ as a weighted sum of its neighbors’ features: $\mathbf{h}_{i}^{\prime\prime}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}W\mathbf{h}_{j}\right)$ where $\sigma$ is a non-linear activation function.
To stabilize the learning process, GAT employs multi-head attention, where $K$ independent attention mechanisms compute the attention coefficients and aggregate the features. The outputs from the different attention heads are concatenated as: $\mathbf{h}_{i}^{\text{multi}}=\big{\|}_{k=1}^{K}\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}^{k}W^{k}\mathbf{h}_{j}\right).$ For the final prediction layer, the attention heads are averaged rather than concatenated: $\mathbf{h}_{i}^{\text{final}}=\sigma\left(\frac{1}{K}\sum_{k=1}^{K}\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}^{k}W^{k}\mathbf{h}_{j}\right).$
In GAT architecture, they use $K=8$ attention heads, with each head computing $F^{\prime}=8$ features, leading to a concatenated output of size $64$ for each node, which results in protein’s final representation size of [1, node_size, 64]. Then we simply perform average pooling to remove the ‘node_size’ dimension. This process yields the latent representation of the protein with size of [1, 64], which will be stored and used later.
A.5 Popular and Rare Protein
For the protein perspective analysis, we randomly selected 10 popular proteins and 10 rare proteins to study. The IDs of the popular proteins are: [4NWH, 2RAY, 4I8S, 3E3D, 4H8Y, 4H92, 1B1V, 3FE0, 3PYK, 3S9T], while the IDs of the rare proteins are: [3I1A, 4Q2G, 4S3K, 3GEU, 3OZQ, 3LOV, 1NVI, 3MGC, 4J77, 2GRM]. Details for these proteins are presented in Figure 9.
@@ A.6 GearNet Training
A.7 Number of Projection Head Layer
To address Research Question 5, we investigate whether increasing the number of layers in the projection head improves alignment performance. As shown in Figure 8, we only modify the number of linear layers in the GDM’s projection head, while leaving the LLM’s projection head unchanged. Our multi-layer projection heads consist of either 2 or 3 linear layers, with ReLU activation functions between each layer.
Figure 8: a. Transition from Single-Layer to Multi-Layer Projection Head. The figure illustrates the shift from a single-layer to a multi-layer projection head architecture. Note, layers are only added to the GDM projection head, while the LLM projection head remains unchanged. b. Number of Projection Head Layer Analysis This extends the example from Figure 5, showing the alignment performance when additional layers are added to the projection head for model pairs involving GAT.
Since each GDM has a different representation dimension, and the output dimension depends on the LLM used in the model pair, we selected different hidden layer dimensions for each GDM’s projection head to ensure appropriate alignment.
@@ A.7 Number of Projection Head Layer
A.8 LLM Finetune Implementation Details
To answer Research Question 6, we fine-tune all three of the LLMs, Gemma2-2B, LLaMa3.1-8B, LLaMa3.1-70B on the protein dataset we constructed during the FASTA preprocessing using LoRA (Low-Rank Adaptation) [9] to efficiently adjust the model’s parameters.
First, the LLM and its corresponding tokenizer were loaded using HuggingFace. The model was loaded with 4-bit quantization to enable efficient memory usage while preserving the model’s performance. Quantization techniques, such as nf4 (normal float 4-bit) and float16 computation, were employed to reduce the computational load during training. The model’s configurations were adjusted to disable caching and to use a single pretraining tensor parallelism thread.
The input data for fine-tuning was formated like this:
@@ A.8 LLM Finetune Implementation Details
For the fine-tuning process, LoRA was applied to enable efficient parameter adjustment without the need to fully fine-tune the LLM. The LoRA configuration specified a low-rank parameter r=8, with lora_alpha=16 and a dropout of 0.05, targeting causal language modeling tasks.
@@
The training arguments were set to use the paged AdamW optimizer in 32-bit mode, with a learning rate of 2e-4 and a cosine learning rate scheduler. The training ran for 10 epoch with a batch size of 1 per device and gradient accumulation steps of 16 to account for memory constraints. Mixed precision (FP16) was used to further optimize memory usage and speed.
We used four A100 GPUs and one A6000 GPU for hardware support. The A100 GPUs were used to fine-tune LLaMa3.1-70B, while the A6000 GPU was sufficient for the remaining experiments.
@@
Figure 9: Protein Examples. The figure provides details of 10 popular proteins and 10 rare proteins. The first column lists the protein IDs, the second column shows the 3D structures of the proteins, and the third column contains the text descriptions for each protein.

### v2

@@ Acknowledgments
Author Contributions
Conceptualization, D.S. and M.D.; Methodology, D.S. and M.D.; Software, D.S.; Investigation, D.S.; Resources, B.D. and M.D.; Writing – Original Draft, D.S., and M.D.; Writing – Review & Editing, All authors; Supervision: K.G., K.Z., J.T., and M.D.
Declaration of Interests
The authors declare no competing interests.
Declaration of Generative AI and AI-assisted technologies in the writing process
The authors used ChatGPT in order to check grammar mistakes.
Supplemental Information
1 Dataset Statistic
2 Prompt Used in FASTA Preprocess
3 Protein Examples
4 GDM Representation Details
5 Popular and Rare Protein
6 GearNet Training
7 Number of Projection Head Layer
8 LLM Finetune Implementation Details
9 Protein-focused MLLMs Methodology
10 Reweighting Methodology
11 Retrieval Methodology.
Figure S1 Data Statistic.
Figure S2 Single-Layer to Multi-Layer Projection Head: Architecture and Results.
Figure S3 Architecture of Protein-Focused MLLMs and Their Generated Responses.
Figure S4 Popular Protein Examples.
Figure S5 Rare Protein Examples.
Figure S6 Example of Protein-focused MLLMs Generated Output for Protein ID 3I1A.
Supplemental Notes
1 Dataset Statistic
Our dataset consists of 20,000 proteins, each with an associated FASTA file and PDB file. Detailed statistics for the dataset are presented in Figure S5, where each pie chart represents a different categorical breakdown of the same dataset. For the “Number of Chains” statistic, the chart shows the count of single-chain proteins versus multiple-chain proteins, with further subdivisions indicating the exact number of chains for multi-chain proteins. For the “Sequence Length” statistic, proteins are categorized by sequence length range. Each section of the pie chart is labeled with the relevant number of chains or sequence length range at the top, with the exact count and corresponding percentage displayed below each label.
2 Prompt Used in FASTA Preprocess
@@
Your task is to summarize complex protein sequence data into two or
three sentences that highlight key features such as molecute type,
chains, structural motifs, organism, etc.
@@
3 Protein Examples
As shown in Figure S8 and S9, we listed several proteins and their corresponding description. The meaning of ‘popular’ and ‘rare’ protein is discussed in Research Question 3: Protein Perspective Analysis.
4 GDM Representation Details
4.1 GearNet
@@
where $f_{i}$ is the initial feature of node $i$, $\mathcal{N}_{r}(i)$ is the set of neighbors connected to node $i$ via edges of type $r$, $W_{r}$ is a learnable weight matrix for edge type $r$, and $\sigma$ is a ReLU activation function. Batch normalization (BN) is applied for normalization across batches, and a residual connection is added from the previous layer to stabilize training. This relational message passing allows GearNet to effectively capture both sequential and spatial information in the protein structure, as different edge types are modeled with different convolutional kernels.
@@
where $m^{(l)}_{(i,j,r)}$ represents the message for edge $(i,j,r)$ at layer $l$, $\mathcal{N}^{\prime}_{r^{\prime}}((i,j,r))$ is the set of neighboring edges in $G^{\prime}$, and $W^{\prime}_{r^{\prime}}$ is a learnable weight matrix for edge type $r^{\prime}$. The angular information between edges is used to determine edge types, with edges having smaller angles expected to interact more strongly. The output from the edge message passing is integrated into the node update step by modifying the aggregation function as follows:
@@
Figure S5: Data Statistic. The pie chart on the left categorizes proteins by the number of chains, while the pie chart on the right categorizes them by sequence length.
4.2 GVP
@@
4.3 ScanNet
Parsing the PDB File:
First, the PDB file is parsed to extract the protein’s amino acid sequence and atomic point cloud. Each atom is represented as a triplet:
@@ 4.3 ScanNet
Local Frame Construction:
Next, a local reference frame is defined for each heavy atom based on the molecular graph of the protein (i.e., nodes represent atoms and edges represent covalent bonds). For each atom, we select two neighboring atoms to construct a local coordinate system. The center is defined by the atom itself, and the orientation is determined by its neighbors using the following equations:
@@
Atomic and Amino Acid Pooling:
ScanNet constructs atomic representations by extracting the local neighborhood of each atom and encoding it with spatio-chemical filters. These filters operate on the atomic coordinates and attributes (such as atom type) and produce an atomic-scale representation. This process is mathematically formalized using spatio-chemical Gaussian filters:
@@
where $G(\mu_{g},\Sigma_{g},x_{k})$ is the Gaussian kernel parameterized by mean $\mu_{g}$ and covariance $\Sigma_{g}$, applied to the local coordinates $x_{k}$ of the neighborhood. The atomic-scale representations are then pooled at the amino acid level. The pooling step aggregates the atom-wise features into a single representation for each amino acid using a multi-head attention mechanism:
@@
Amino Acid Neighborhood Embedding and Protein Representation:
For each amino acid, its local neighborhood is constructed based on the $C_{\alpha}$ atom, sidechain orientation, and backbone orientation. The same spatio-chemical filtering process is applied to obtain an amino acid-wise representation:
@@
4.4 GAT
For a graph with $N$ nodes, the input to the GAT layer is a set of node features $\mathbf{h}=\{\mathbf{h}_{1},\mathbf{h}_{2},\dots,\mathbf{h}_{N}\}$, where $\mathbf{h}_{i}\in\mathbb{R}^{F}$ is the feature vector of node $i$, and $F$ is the number of features associated with each atom. Each node is first transformed linearly using a shared learnable weight matrix $W\in\mathbb{R}^{F^{\prime}\times F}$, yielding: $\mathbf{h}_{i}^{\prime}=W\mathbf{h}_{i}$. Then, the attention mechanism computes attention coefficients for each pair of nodes $i$ and $j$, where $j\in\mathcal{N}_{i}$ is a neighboring node of $i$. The attention coefficients $e_{ij}$ are given by: $e_{ij}=\text{LeakyReLU}\left(\mathbf{a}^{T}\left[W\mathbf{h}_{i}\parallel W\mathbf{h}_{j}\right]\right)$, where $\mathbf{a}\in\mathbb{R}^{2F^{\prime}}$ is a learnable weight vector, and $\parallel$ denotes concatenation. These attention coefficients are then normalized across all neighbors of node $i$ using the softmax function: $\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}.$ The normalized attention coefficients are used to compute the updated feature representation of node $i$ as a weighted sum of its neighbors’ features: $\mathbf{h}_{i}^{\prime\prime}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}W\mathbf{h}_{j}\right)$ where $\sigma$ is a non-linear activation function.
To stabilize the learning process, GAT employs multi-head attention, where $K$ independent attention mechanisms compute the attention coefficients and aggregate the features. The outputs from the different attention heads are concatenated as: $\mathbf{h}_{i}^{\text{multi}}=\big{\|}_{k=1}^{K}\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}^{k}W^{k}\mathbf{h}_{j}\right).$ For the final prediction layer, the attention heads are averaged rather than concatenated: $\mathbf{h}_{i}^{\text{final}}=\sigma\left(\frac{1}{K}\sum_{k=1}^{K}\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}^{k}W^{k}\mathbf{h}_{j}\right).$ In GAT architecture, they use $K=8$ attention heads, with each head computing $F^{\prime}=8$ features, leading to a concatenated output of size $64$ for each node, which results in protein’s final representation size of [1, node_size, 64]. Then we simply perform average pooling to remove the ‘node_size’ dimension. This process yields the latent representation of the protein with size of [1, 64], which will be stored and used later.
5 Popular and Rare Protein
For the protein perspective analysis, we randomly selected 10 popular proteins and 10 rare proteins to study. The IDs of the popular proteins are: [4NWH, 2RAY, 4I8S, 3E3D, 4H8Y, 4H92, 1B1V, 3FE0, 3PYK, 3S9T], while the IDs of the rare proteins are: [3I1A, 4Q2G, 4S3K, 3GEU, 3OZQ, 3LOV, 1NVI, 3MGC, 4J77, 2GRM]. Details for these proteins are presented in Figure S8 and S9.
@@ 6 GearNet Training
7 Number of Projection Head Layer
To address Research Question 5, we investigate whether increasing the number of layers in the projection head improves alignment performance. As shown in Figure S6A, we only modify the number of linear layers in the GDM’s projection head, while leaving the LLM’s projection head unchanged. Our multi-layer projection heads consist of either 2 or 3 linear layers, with ReLU activation functions between each layer. Since each GDM has a different representation dimension, and the output dimension depends on the LLM used in the model pair, we selected different hidden layer dimensions for each GDM’s projection head to ensure appropriate alignment.
Figure S6: Single-Layer to Multi-Layer Projection Head: Architecture and Results. (A) The figure illustrates the shift from a single-layer to a multi-layer projection head architecture. Note, layers are only added to the GDM projection head, while the LLM projection head remains unchanged. (B) This extends the example from Figure 4A, showing the alignment performance when additional layers are added to the projection head for model pairs involving GAT.
@@ 7 Number of Projection Head Layer
8 LLM Finetune Implementation Details
To answer Research Question 6, we fine-tune all three of the LLMs, Gemma2-2B, LLaMa3.1-8B, LLaMa3.1-70B on the protein dataset we constructed during the FASTA preprocessing using LoRA (Low-Rank Adaptation) 38 to efficiently adjust the model’s parameters.
First, the LLM and its corresponding tokenizer were loaded using HuggingFace. The model was loaded with 4-bit quantization to enable efficient memory usage while preserving the model’s performance. Quantization techniques, such as nf4 (normal float 4-bit) and float16 computation, were employed to reduce the computational load during training. The model’s configurations were adjusted to disable caching and to use a single pretraining tensor parallelism thread. The input data for fine-tuning was formated like this:
@@ 8 LLM Finetune Implementation Details
For the fine-tuning process, LoRA was applied to enable efficient parameter adjustment without the need to fully fine-tune the LLM. The LoRA configuration specified a low-rank parameter r=8, with lora_alpha=16 and a dropout of 0.05, targeting causal language modeling tasks. The training arguments were set to use the paged AdamW optimizer in 32-bit mode, with a learning rate of 2e-4 and a cosine learning rate scheduler. The training ran for 10 epoch with a batch size of 1 per device and gradient accumulation steps of 16 to account for memory constraints. Mixed precision (FP16) was used to further optimize memory usage and speed. We used four A100 GPUs and one A6000 GPU for hardware support. The A100 GPUs were used to fine-tune LLaMa3.1-70B, while the A6000 GPU was sufficient for the remaining experiments.
9 Protein-focused MLLMs Methodology
Figure S7: Architecture of Protein-Focused MLLMs and Their Generated Responses. (A) The architecture of a protein-focused MLLM comprises three primary components: a Geometric Deep Model (GDM), a projection head, and a Large Language Model (LLM). The GDM processes the 3D protein structures as its input. Subsequently, the projection head transforms these structural representations into the same dimension as the LLM. Then LLM takes both projected representation and textual inputs to generate its outputs. (B) The figure shows protein-focused MLLMs example responses for protein ID 4NWH. In the output, factually correct text is highlighted in blue, whereas text that is incorrect or involves hallucination is marked in yellow.
As shown in Figure S7, our protein-focused MLLM contains three key components: a GDM, a trained projection head, and a LLM. The protein-focused MLLM can take two inputs: a protein structure and a text question. In this study, we focus on the protein description task, where the text input is “Describe the protein in detail.” First, the protein structure is processed by the GDM to extract a latent representation. This representation is fed into the trained projection head, which maps it to the same dimensional space as the LLM’s embeddings. The text input is tokenized and passed through the LLM to extract its representation. We then concatenate the projected protein representation with the text representation and feed them into the LLM to generate the output. This output will be used when we further fine-tuned the projection head to improve alignment between the protein and text representations. We use the protein descriptions that generated during the FASTA preprocessing section as the ground truth. We fine-tuned the projection head using Cross-Entropy Loss, which calculates the discrepancy between the predicted token sequence and the ground truth sequence:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\log P\left(y_{t}^{(i)}\mid x^{(i)},y_{<t}^{(i)}\right)$
where $N$ is the number of samples in the batch. $T$ is the length of the ground truth description for the $i$-th sample. $y_{t}^{(i)}$ is the ground truth token at position $t$ for the $i$-th sample. $x^{(i)}$ is the input to the model. $y_{<t}^{(i)}$ is all ground truth tokens generated prior to $t$ for the $i$-th sample. During fine-tuning, the GDM and LLM parameters were frozen, and only the projection head was updated. We used the same implementation details as in the Representation Alignment section. The projection head was fine-tuned for 40 epochs with learning rate of $1\times 10^{-3}$ and batch size of 32.
@@ 9 Protein-focused MLLMs Methodology
Figure S8: Popular Protein Examples. The figure provides details of 10 popular proteins. The first column lists the protein IDs, the second column shows the 3D structures of the proteins, and the third column contains the text descriptions for each protein.
Figure S9: Rare Protein Examples. The figure provides details of 10 rare proteins. The first column lists the protein IDs, the second column shows the 3D structures of the proteins, and the third column contains the text descriptions for each protein.
10 Reweighting Methodology
We reweight the penalty of the rare protein in the loss function when training the projection head. The process starts by labeling protein as rare and popular. We used the same way as discussed in Supplemental Note 5 to rank top 100 rare protein from the training dataset. During the training, we reweight the penalty of these 100 rare protein in the loss function, by multiplying rare protein penalties by 2, and leave other protein penalties unchanged. If protein $i$ in 100 rare protein:
$\mathcal{L}_{\text{total}}=\frac{1}{B}\sum_{i=1}^{B}\left(-\log\left(\frac{\exp\left(\frac{\text{sim}(g_{i},t_{i})+1}{2\tau}\right)}{\exp\left(\frac{\text{sim}(g_{i},t_{i})+1}{2\tau}\right)+\sum_{j\neq i}\exp\left(\frac{\text{sim}(g_{i},t_{j})+1}{2\tau}\right)}\right)*2\right)$
11 Retrieval Methodology
The retrieval process begins by mapping all training protein representations to the same dimensional space as the LLM using our reweighting-trained projection head. This mapped representation space is denoted as $R$. Once all protein representations in the training dataset are mapped, each protein $j$ in the testing set is also mapped into $R$. For each protein $j$, we identify its top $k$ most similar proteins in the representation space $R$ based on the cosine similarity with other protein $i$:
[…]
In this study, we evaluate $k$ with different values: $[3,5,10]$. For each test protein $j$, the ground truth descriptions of its top $k$ similar proteins are prepended to the original input and fed into the protein-focused MLLM. The model’s performance is then evaluated using the ROUGE and BLEU metrics.
Figure S10: Example of Protein-focused MLLMs Generated Output for Protein ID 3I1A. The upper figure illustrates the output generated by the model using only the reweighting technique. The lower figure, on the other hand, displays the output generated by the model when both reweighting and retrieval techniques are employed.